OpenStudy (anonymous):

cos(x+y) + cos(x+y) = 2cos(x+y) What is this process called again? Simplification? Sorry, having a slow moment.

OpenStudy (kinggeorge):

Simplification would be a good term for that.

OpenStudy (anonymous):

http://ocw.mit.edu/resources/res-18-005-highlights-of-calculus-spring-2010/derivatives/derivative-of-sin-x-and-cos-x/ The two key functions of oscillation have specially neat derivatives: The slope of sin x is cos x ! And the slope of cos x is - sin x. These come from one crucial fact: (sin x) / x approaches 1 at x = 0. This checks that the slope of sin x is cos 0 = 1 at the all-important point x = 0. Professor Strang connects sine and cosine to moving around a circle, or up and down for a spring, or in and out for your lungs. Professor Strang's Calculus textbook (1st edition, 1991) is freely available here. Subtitles are provided through the generous assistance of Jimmy Ren.

hero (hero):

How about a more obvious explanation: x + x = 2x 2x = 2x x^2 + x^2 = 2x^2 2x^2 = 2x^2 (a+b) + (a+b) = 2(a+b) 2(a+b) = 2(a+b) cos(x+y) + cos(x+y) = 2cos(x+y) 2cos(x+y) = 2cos(x+y)

hero (hero):

It's called adding like terms by the way